OpenStudy (anonymous):

1. Let f(x)=x^3-2x. Simplify the expression f(x+h)-f(x)/h 2. Let f(x)=3x+2 and g(x)=x/x+1 3. Let f(x)=2x^2. Simplify the expression f(x+h)-f(x)/h

OpenStudy (anonymous):

wherever there is an f(x) in the equation to the right, plug in x^3-2x

OpenStudy (akashdeepdeb):

Here's something that might help, since you're asking similar questions. 1. \(f(x) = x^3-2x\) This is a function. Basically, this means that: \(f(a) = a^3-2x\) \(f(b) = b^3 - 2x\) \(f(...) = (...)^3 - 2(...)\) \(f('conad') = ('conad')^3 - 2('conad')\) \(f(x+h) = (x+h)^3 - 2(x+h)\) Now, you just have to find: \[\frac{f(x+h) - f(x)}{h}\] Try it.